OpenStudy (anonymous):

A number added to its square is 56. Find the number. @noemy_bieber

OpenStudy (anonymous):

7 Or -8

OpenStudy (anonymous):

Let the number be x x+x^2=56 x^2+x-56=0 x^2+8x-7x-56=0 (x-7)(x+8)=0 Either 7 or -8